# Parameter estimation of microlensed gravitational waves with Conditional Variational Autoencoders

### v1

@@ VI Conclusions and outlook
We have studied the suitability of Conditional Variational Autoencoders (CVAE) to conduct parameter inference of microlensed gravitational waves from binary black hole mergers. Specifically, we have trained a CVAE to estimate the lensing parameters of a point mass lens model, namely $y$ (the relative source position) and $t_{M}$ (the characteristic lensing time). Our study has shown that CVAE are quite capable to perform this task, yielding high performance in terms of both accuracy and computational efficiency.
The CVAE model has displayed substantial advantages over traditional Bayesian methods, including a significant reduction in computation time, up to five orders of magnitude faster than bilby. This efficiency makes CVAE an attractive option for real-time or low-latency applications in GW data analysis, addressing one of the main challenges in current microlensing studies, computational cost. Furthermore, our analysis has revealed that incorporating CVAE-generated priors in the inference can further enhance Bayesian frameworks like bilby. This suggests a promising hybrid approach that combines the rapidness of machine learning and the reliability of traditional statistical inference.
Future research can focus on refining the architecture of the CVAE and making it more suitable for real-world applications. First and foremost, it will be essential to simulate the data using detector noise. In addition, the inclusion of normalizing flows, as proposed by [76], may enhance the model’s performance. Furthermore, incorporating spectrograms as extra input data could enable the CVAE to extract additional features, potentially leading to more precise parameter estimates. Finally, trying the model on real data will be an interesting exercise that could potentially bring new insights to the field of gravitational lensing.
@@ Acknowledgements
The authors are grateful to Mick Wright and Sreekanth Harikumar for their valuable support and guidance with the Gravelamps package. We also thank Narenraju Nagarajan and Melissa Lopez Portilla for their helpful comments on the manuscript. This work was supported by the Spanish Ministry of Science and Innovation (grants PID2021-125485NB-C21 and PID2021-125485NB-C22 funded by MCIN/AEI/10.13039/501100011033 and ERDF, ‘A way of making Europe’), along with the Agència de Gestió d’Ajuts Universitaris i de Recerca (AGAUR), Generalitat de Catalunya (grant SGR-2021-01069), and the Generalitat Valenciana (grant CIPROM/2022/49). RBN acknowledges financial support from NTT DATA Spain and expresses gratitude to Prof. Javier Béjar for his guidance on machine learning matters. OGF is supported by an FCT doctoral scholarship (reference UI/BD/154358/2022). Computations have been performed at the Nyx cluster of the ICCUB. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.
@@ Appendix A Features of the CVAE model
since we use no padding or dilatation. All the layers, except for the last linear layer of the encoder, decoder and recognition networks, have Leaky ReLU as activation function. The weights of the convolutional block are initialized using the He method at the beginning of the training process. With respect to the CVAE presented in [75], minor changes were carried out, most of them concerning the dimensions of the layers. This was motivated by the fact that our input signal is longer and we are estimating only two parameters. We also added an extra convolutional layer in the shared convolutional block. These decisions were taken in part by trial and error, and the final configuration is the one that achieved the best results.
@@
Once the CVAE is trained we can test it on a test set. We use a set of 250 waveforms that have never been seen by the model before to test its generalization capability. During the testing process, the recognition encoder is not used. Instead, the time series is passed through the convolutional block and then through the encoder. A sample is taken from the latent space, concatenated with the output of the convolutional block and passed through the decoder, obtaining the parameters of two truncated Gaussian distributions. By passing the same time series $N$ times through the model, one gets $N$ different pairs of truncated gaussians. We take a sample from each of these distributions, obtaining a set of $N$ samples that define a posterior probability that can be represented in a corner plot. When $N=8000$ this process takes, on average, around 4 seconds per waveform. The GPU used for training and testing is an NVIDIA A40.

### v2

@@ VI Conclusions and outlook
We have studied the suitability of Conditional Variational Autoencoders (CVAE) to conduct parameter inference of microlensed GWs from binary black hole mergers. Specifically, we have trained a CVAE to estimate the lensing parameters of a point mass lens model, namely $y$ (the relative source position) and $t_{M}$ (the characteristic lensing time). Our study has shown that CVAE are quite capable to perform this task, yielding high performance in terms of computational efficiency.
The CVAE model has displayed substantial advantages over traditional Bayesian methods in computation time, up to five orders of magnitude faster than Bilby, at the expense of some loss in accuracy. This efficiency makes CVAE an attractive option for real-time or low-latency applications in GW data analysis, addressing one of the main challenges in current microlensing studies, computational cost. Furthermore, our analysis has revealed that incorporating CVAE-generated priors in the inference can further enhance Bayesian frameworks like Bilby. This suggests a promising hybrid approach that combines the rapidness of machine learning and the reliability of traditional statistical inference.
Future research can focus on refining the architecture of the CVAE and making it more suitable for real-world applications. First and foremost, it will be essential to simulate the data using detector noise. In addition, the inclusion of normalizing flows, as proposed by [83], may enhance the model’s performance. Normalizing flows have proven to be a very powerful tool for parameter estimation in GW science [85, 86, 83], and hence they can be an important way of improving the model. Furthermore, incorporating spectrograms as extra input data could enable the CVAE to extract additional features present in the time-frequency representation, potentially leading to more precise parameter estimates. Enlarging the number of parameters estimated by the CVAE is also a future goal. Finally, applying the model to real data could serve as a valuable validation step.
@@ Acknowledgements
The authors are grateful to Mick Wright and Sreekanth Harikumar for their valuable support and guidance with the Gravelamps package. We also appreciate Narenraju Nagarajan and Melissa Lopez Portilla for their helpful comments on the manuscript. Finally, we thank the anonymous referee for their useful and constructive feedback on the first version of this manuscript. This work was supported by the Spanish Ministry of Science and Innovation (grants PID2021-125485NB-C21 and PID2021-125485NB-C22 funded by MCIN/AEI/10.13039/501100011033 and ERDF, ‘A way of making Europe’), along with the Agència de Gestió d’Ajuts Universitaris i de Recerca (AGAUR), Generalitat de Catalunya (grant SGR-2021-01069), and the Generalitat Valenciana (grant CIPROM/2022/49). RBN acknowledges financial support from NTT DATA Spain and expresses gratitude to Prof. Javier Béjar for his guidance on machine learning matters. OGF is supported by an FCT doctoral scholarship (reference UI/BD/154358/2022). Computations have been performed at the Nyx cluster of the ICCUB. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.
@@ Appendix A Features of the CVAE model
since we use no padding or dilatation. All the layers, except for the last linear layer of the encoder, decoder and recognition networks, have Leaky ReLU [118] as activation function. The weights of the convolutional block are initialized using the He method at the beginning of the training process. With respect to the CVAE presented in [82], minor changes were carried out, most of them concerning the dimensions of the layers. This was motivated by the fact that our input signal is longer and we are estimating only two parameters. We also added an extra convolutional layer in the shared convolutional block. These decisions were taken in part by trial and error, and the final configuration is the one that achieved the best results.
@@
Once the CVAE is trained we can test it on a test set. We use a set of 1000 waveforms that have never been seen by the model before to test its generalization capability. During the testing process, the recognition encoder is not used. Instead, the time series is passed through the convolutional block and then through the encoder. A sample is taken from the latent space, concatenated with the output of the convolutional block and passed through the decoder, obtaining the parameters of two truncated Gaussian distributions. By passing the same time series $N$ times through the model, one gets $N$ different pairs of truncated gaussians. We take a sample from each of these distributions, obtaining a set of $N$ samples that define a posterior probability that can be represented in a corner plot. When $N=8000$ this process takes, on average, around 4 seconds per waveform. The GPU used for training and testing is an NVIDIA A40.